# Alberta Wells Dataset: Pinpointing Oil and Gas Wells from Satellite Imagery

### v2

@@ 4 Experiments
We augment images by randomly resizing images to 256$\times$256, ensuring all bounding boxes remain intact for object detection. We then apply horizontal and vertical flipping with a probability of 0.25 each, followed by normalization using channel-wise mean and standard deviation calculated from the training split of the dataset.
The hyperparameters we use in these various models represent standard performant settings and are not intended to represent the outcome of hyperparameter optimization.
@@ 4.1 Binary Segmentation
U-Net [33] was chosen for its widespread use as a baseline, offering an effective encoder-decoder architecture for multi-scale feature extraction. PAN [34] improves multi-scale context with pyramid pooling and attention mechanisms. DeepLabV3+ [35] was selected for its popularity in remote sensing tasks with its Atrous Convolution and ASPP module for capturing contextual information at various scales. SegFormer [36] is a transformer-based architecture designed for semantic segmentation, utilizing self-attention mechanisms for capturing long-range dependencies. UperNet [37] combines UNet and PSPNet [38] architectures, featuring a UNet-like structure for multi-scale feature fusion and PSPNet’s pyramid pooling module integrated with a Swin Transformer [39] backbone for efficient multi-scale processing.
We train all CNN-based models using a ResNet50 backbone, a batch size of 128, and the BCELogits loss function. To fine-tune the model, a cosine annealing scheduler [40] is used, which adjusts the learning rate smoothly in a cyclical manner by gradually decreasing it. We also experimented with additional backbones with U-Net to evaluate the impact of backbones with larger receptive fields and attention mechanisms. This included ResNeXt50 [41], which enhances feature learning through grouped convolutions; SE-ResNet50 [42], which introduces channel-wise attention with Squeeze-and-Excitation blocks; and EfficientNetB6 [43], known for its balanced scaling.
In segmentation, the combination of UperNet’s contextual aggregation and the ConvNeXt backbone’s efficient representation enables precise segmentation, particularly valuable for remote sensing applications. For transformer-based models, while both Segformer and UperNet use a Dice loss function and a polynomial learning rate scheduler, Segformer utilizes a mit-b0-ade [36] backbone with a batch size of 128, while UperNet uses ConvNeXT (small and base) and Swin Transformer backbones with a batch size of 64. All models are optimized using AdamW for 50 epochs.
4.2 Object Detection
For binary object detection, we consider both single-stage, i.e., RetinaNet [44], FCOS and SSD, and two-stage CNN-based architectures, i.e. Faster R-CNN [45].
RetinaNet [44] is a one-stage architecture trained using focal loss, which helps to address class imbalance. It uses a Feature Pyramid Network (FPN) for multi-scale feature extraction and efficient object detection across different scales. Faster R-CNN [45] is a two-stage model recognized for its high accuracy. It employs a Region Proposal Network (RPN) for generating region proposals and a separate network for predicting class labels and refining bounding box coordinates. FCOS (Fully Convolutional One-Stage Object Detection) [46] directly predicts object locations and categories from feature maps, which is effective for small object detection. SSD (Single Shot MultiBox Detector) [47] uses multiple feature maps at different scales, enhancing its accuracy for small objects.
All object detection models are trained using a ResNet50 backbone, except for SSD Lite, which is trained with a MobileNet backbone. The batch size is set to 256 for Faster R-CNN and FCOS and 512 for RetinaNet and SSD Lite. We used a cosine annealing scheduler [40] and trained all models for 120 epochs. All models are optimized using the AdamW optimizer.
For binary object detection model evaluation, we calculate Intersection over Union (IoU) at various thresholds (e.g., IoU0.1, IoU0.3, IoU0.5), which measures how well predicted bounding boxes align with ground truth. IoU is computed by dividing the area of overlap by the area of their union, with higher values indicating better alignment. IoU thresholds define the minimum overlap required for a predicted box to match a ground truth box. (For example, an IoU0.5 threshold means a predicted box must have at least 50% overlap with a ground truth box to be considered a correct detection.)
We also assess Mean Average Precision (mAP), including mAP50 and mAP50:95, measuring the model’s precision-recall trade-off and detection accuracy at various IoU thresholds. mAP50 measures precision at an IoU threshold of 0.5, while mAP50:95 averages precision across IoU thresholds from 0.5 to 0.95. Higher mAP scores reflect better detection accuracy and precision.While higher IoU values indicate better accuracy for individual predictions, mAP provides a broader measure of detection performance by capturing precision across different IoU criteria.
Figure 7: A sample image patch from our dataset includes examples with no wells, two wells, and multiple wells. Additionally, we present qualitative results with predictions generated by our Segmentation U-Net (EfficientNet-B6) and Detection FCOS model.
5 Results
Our tasks involve identifying a roughly circular well region with a 90m diameter in real life, which translates to less than 30 pixels in satellite imagery due to resizing and other augmentations.
This poses a challenge for machine learning models given the heterogeneous nature of the background, including various similarly shaped and sized features of the natural and built environment. Additionally, vegetation can occlude wells in RGB channels, highlighting the importance of near-infrared imagery for guidance. The wells themselves also vary somewhat in shape and can be in various states of disrepair as a result of differing ages and maintenance.
5.1 Binary Segmentation
For the binary segmentation task framing, we train Models (from scratch) using both CNN-based and Transformer-based backbones, considering the prevalent imbalance in the image data due to the small size of wells. Although we did use 3-dimensional, ImageNet initialized weights of the backbone but modified the initial layers afterwards to support 4-dimensional multispectral images.
Among our models, as shown in Table 4, the traditional U-Net with EfficientNetB6 backbone performs the best, with CNN-based models showing the highest IOU of $60.4\pm 0.3$ and F1-Score of $64.8\pm 0.4$. While a ResNet50-based backbone achieves the highest Precision of $90.2\pm 2.2$, indicating more accurate predictions of well instances compared to other models. Precision, which reflects the accuracy of our positive detections compared to the ground truth, is crucial. However, a high recall value ensures the model captures most actual well instances, reducing the risk of missing important information. Thus, the Uper-Net model with the highest recall value of $73.1\pm 0.1$, which excels at capturing global context information, appears a good candidate for this task, given a decent precision score.
Taking into account both precision and recall, U-Net with EfficientNetB6 backbone performs well, suggesting the utility of a larger backbone with a bigger receptive field. However, ConvNeXt Small provides a trade-off between computational efficiency and performance, requiring fewer GFLOPs and parameters compared to UperNet with Swin Small while achieving similar performance.
@@ 5.2 Binary Object Detection
Table 6: Results for the binary segmentation task for U-Net with ResNet50 backbone, trained with and without NIR imagery. We find that NIR significantly improves performance.
@@
Table 7: Results for object detection task for the FCOS Model with ResNet50 backbone evaluated over the test set for multiple input modality. We report the intersection over union (IoU) over thresholds $0.1,0.3,0.5$ and the mean average precision (mAP) for both IoU$=0.5$ and IoU$\in[0.5,0.95]$ thresholds.
Modality IoU0.1 IoU0.3 IoU0.5 mAP50 mAP50:95
@@
Our evaluation, as shown in Table 5, indicates that while all models perform reasonably well in terms of aligning predicted and actual well locations, performance in the object detection task is overall lower than for segmentation – indicating that potentially segmentation is simply a better framing for this task in real-world settings.
The observed gap in performance is likely due to the small size of the wells. It is well-known that single-stage CNN architectures (such as FCOS and SSD) often demonstrate better performance on small object detection than two-stage methods (such as Faster R-CNN), and this aligns with our observations. The exception here is the single-stage method RetinaNet, which, although it has comparable IoU scores, struggles to detect wells accurately.
SSD Lite stands out with the highest IoU0.5 score of $65.07\pm 0.03$ and IoU0.3 score of $50.3\pm 0.08$. Whereas all models are quite similar in terms of IoU0.1, the highest score by FasterRCNN is $36.79\pm 1.07$. Thus, SSD Lite and FCOS excel in localization, especially at higher IoU thresholds, while Faster R-CNN is adept at detecting objects with minimal overlap. All models demonstrate low performance in terms of mAP50, which assesses precision-recall trade-off and detection accuracy at an IoU threshold of 0.5. FCOS achieves the of $9.67\pm 1.47$ while SSD Lite achieves a score of $9.76\pm 0.39$. This may be due to these models not producing region proposals confidently enough, especially in instances with a large number of wells. Whereas over a broader evaluation with mAP50:95 which averages precision across IoU thresholds from 0.5 to 0.95. All models apart from RetinaNet provide much better results, with FCOS achieving the highest score of $30.46\pm 3.11$, indicating a decent performance in the identification of well instances. While IoU scores are decent with increasing thresholds, the mAP scores indicate that a more complex model may be required for this task.
5.3 Value Added from Dataset Breadth
5.3.1 Benefits of NIR imagery: (RGB v/s RGB+NIR)
We find that including the NIR band significantly enhances both object detection and segmentation performance compared to purely RGB data. In object detection (Table 7), RGB+NIR achieves higher Intersection over Union (IoU) scores across all thresholds (0.1, 0.3, and 0.5) and a considerable increase in mAP@50 (9.67$\pm$1.47 vs. 5.7$\pm$3.65) and mAP@50:95 (30.46$\pm$3.11 vs. 20$\pm$10.4). Similarly, in segmentation (Table 6), RGB+NIR shows superior performance in IoU (58$\pm$0.5 vs. 56.6$\pm$0.44), F1 Score (61.9$\pm$0.8 vs. 60.5$\pm$0.35), and Precision (90.2$\pm$2.2 vs. 87$\pm$1.4), while maintaining a slightly lower Recall (62.3$\pm$1.6 vs. 62.54$\pm$0.13). These improvements are due to the enhanced spectral information, which is particularly effective in identifying features such as ground depressions, often more visible in the NIR band.
5.3.2 Benefits of Using Multiple Well Types.
While much past work has focused on active wells, we show that training on such data is insufficient to detect abandoned and suspended wells (which are often significantly harder to e.g. degradation and vegetation growth over time). The results in Table 8 emphasize the necessity of including all well types (active, suspended, and abandoned) in the training dataset to ensure comprehensive detection across diverse scenarios. While training exclusively on active wells achieves high performance for individual well types, but it significantly under-performs in mixed-type contexts (IoU: 0.502), indicating poor generalization when all well types are present. Training exclusively on active wells significantly under-performs in mixed-type contexts (IoU: 0.502), indicating poor generalization when all well types are present. Conversely, training on all well types improves the model’s ability to handle real-world heterogeneity, as reflected by a higher IoU (0.576), F1 score (0.614), and recall (0.614) for the test set. The enhanced recall demonstrates the model’s capability to identify a broader range of wells, crucial for environmental monitoring, where missing even a single abandoned well could result in unaddressed methane emissions or groundwater contamination. Precision for active-only training (0.998) is higher than the model trained on all well types (0.913), but its inability to detect all wells in an image (due to low recall) limits its applicability. Therefore, incorporating all well types in training ensures balanced, reliable performance, allowing for accurate detection and classification in diverse and realistic contexts.
@@ 6 Limitations
One potential limitation of our work is that we rely on well locations listed by the Alberta Energy Regulator. It is likely that some true well locations are missing in this data, leading to the potential for false negatives in the ground-truth data for this problem. However, it is to be expected that this will not significantly affect the training of algorithms since these labels represent a small fraction of the negative locations in the dataset, and deep learning algorithms are known to be robust to moderate amounts of label noise (see e.g. [48]). Instead the effect may simply be that the reported test accuracy is actually lower than the true value (due to certain correctly predicted well locations being evaluated as false). We hope to investigate such effects further in future work.
@@ 7 Conclusion
In this paper, we present the first large-scale benchmark dataset aimed at identifying oil and gas wells, with a focus on abandoned and suspended wells, which are a significant source of greenhouse gases and other pollutants.
We combine high-resolution imagery, an extensive database of well locations, and expert verification to create the Alberta Wells Dataset. We frame well identification both in terms of object detection and binary segmentation and evaluate the performance of a wide range of popular deep learning methods on these tasks. We find that the UperNet Model (with a ConvNexT backbone), in particular, represents the most promising baseline for the binary segmentation task, while for object detection, all models demonstrate more mixed results, with Single Stage Models (such as FCOS and SSD) providing a relatively promising baseline. These results show that the Alberta Wells Dataset represents both a challenging as well as a societally impactful set of tasks.
The value added by the dataset is twofold. First, most global fossil fuel-producing regions do not have databases of well locations comparable to that provided by AER. Alberta’s varied landscape provides an ideal setting for developing algorithms for the detection of wells, which can then be used directly in other locations or fine-tuned. Second, even in Alberta, the list developed by AER is not comprehensive, and many abandoned wells are believed to be missing; remote sensing algorithms can provide candidate locations for domain experts to examine so as to determine the true number and locations of abandoned wells in Alberta. Overall, we hope that our work will make it possible to (i) better estimate the total effects of well-related methane emissions (currently extremely uncertain [1]) and groundwater contamination, (ii) prioritize wells to be plugged and decontaminated. We believe that the scalability of remote sensing will make it an invaluable tool in mitigating the global environmental impact of abandoned oil and gas wells.
8 Acknowledgments and Disclosure of Funding
This research was made possible by the generous access to the Planet Lab subscription provided by Dr. Tim Elrick from McGill University’s Geography Department. We acknowledge the essential support from Planet Labs for their academic subscription and guidance on the public release of datasets. The Alberta Energy Regulator’s AER ST37 dataset was a crucial resource for our work.
The authors are also grateful to Mila and the Digital Research Alliance of Canada for the vital computing facilities, and to NVIDIA Corporation for their indispensable computational resources. Special thanks to Francis Pelletier, the lab’s Data Engineer, whose crucial feedback on our data collection pipeline and coding significantly improved our methods.
9 Impact Statement
@@
By improving detection techniques and providing access to valuable data, this work has the potential to support global efforts in identifying and addressing abandoned oil and gas wells, ultimately contributing to environmental sustainability and climate action.
Appendix A Data Licensing
The AWD Dataset is released under a Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) License (https://creativecommons.org/licenses/by-nc/4.0/). The satellite imagery for this project was acquired through Planet Labs’ [9] Education & Research license, which allows the use of the data in publications and the creation of derivative products related to those publications. However, the raw imagery cannot be shared publicly. To adhere to these guidelines, we provide the data in HDF5 format, with the satellite imagery pre-processed to produce a derived product represented as a numpy array from Raster Vector. This process removes all geographic metadata.
This data is for academic use only and should not be used commercially. Proper credit to the current authors, Planet Labs [9], and the Alberta Energy Regulator [31] is required when using this data.

### v3

@@ 4 Experiments
We augment images by randomly resizing images to 256$\times$256, ensuring all bounding boxes remain intact for object detection. We then apply horizontal and vertical flipping with a probability of 0.25 each, followed by normalization using channel-wise mean and standard deviation calculated from the training split of the dataset. The hyperparameters we use in these various models represent standard performant settings and are not intended to represent the outcome of hyperparameter optimization.
@@ 4.1 Binary Segmentation
DETR ResNet-50 41.47M 6.86 41.78$\pm$0.11 51.15$\pm$0.14 63.17$\pm$0.11 15.22$\pm$0.28 38.45$\pm$0.32
U-Net (Ronneberger et al., 2015) was chosen for its widespread use as a baseline, offering an effective encoder-decoder architecture for multi-scale feature extraction. PAN (Li et al., 2018) improves multi-scale context with pyramid pooling and attention mechanisms. DeepLabV3+ (Chen et al., 2018) was selected for its popularity in remote sensing tasks with its Atrous Convolution and ASPP module for capturing contextual information at various scales. SegFormer (Xie et al., 2021) is a transformer-based architecture designed for semantic segmentation, utilizing self-attention mechanisms for capturing long-range dependencies. UperNet (Xiao et al., 2018) combines UNet and PSPNet (Zhao et al., 2017) architectures, featuring a UNet-like structure for multi-scale feature fusion and PSPNet’s pyramid pooling module integrated with a Swin Transformer (Liu et al., 2021) backbone for efficient multi-scale processing.
We train all CNN-based models using a ResNet50 backbone, a batch size of 128, and the BCELogits loss function. To fine-tune the model, a cosine annealing scheduler (Loshchilov & Hutter, 2017) is used, which adjusts the learning rate smoothly in a cyclical manner by gradually decreasing it. We also experimented with additional backbones with U-Net to evaluate the impact of backbones with larger receptive fields and attention mechanisms. This included ResNeXt50 (Xie et al., 2017), which enhances feature learning through grouped convolutions; SE-ResNet50 (Hu et al., 2018), which introduces channel-wise attention with Squeeze-and-Excitation blocks; and EfficientNetB6 (Tan & Le, 2019), known for its balanced scaling. In segmentation, the combination of UperNet’s contextual aggregation and the ConvNeXt backbone’s efficient representation enables precise segmentation, particularly valuable for remote sensing applications. For transformer-based models, while both Segformer and UperNet use a Dice loss function and a polynomial learning rate scheduler, Segformer utilizes a mit-b0-ade (Xie et al., 2021) backbone with a batch size of 128, while UperNet uses ConvNeXT (small and base) and Swin Transformer backbones with a batch size of 64. All models are optimized using AdamW for 50 epochs.
4.2 Object Detection
For binary object detection, we consider both single-stage, i.e., RetinaNet (Lin et al., 2017), FCOS and SSD, and two-stage CNN-based architectures, i.e. Faster R-CNN (Ren et al., 2017), and the transformer-based architecture DETR (Carion et al., 2020). RetinaNet (Lin et al., 2017) is a one-stage architecture trained using focal loss, which helps to address class imbalance. It uses a Feature Pyramid Network for multi-scale feature extraction and efficient object detection across different scales. Faster R-CNN (Ren et al., 2017) is a two-stage model recognized for its high accuracy. It employs a Region Proposal Network for generating region proposals and a separate network for predicting class labels and refining bounding box coordinates. FCOS (Fully Convolutional One-Stage Object Detection) (Tian et al., 2019) directly predicts object locations and categories from feature maps, which is effective for small object detection. SSD (Single Shot MultiBox Detector) (Liu et al., 2016) uses multiple feature maps at different scales, enhancing its accuracy for small objects. DETR (DEtection TRansformers) is a transformer-based model that treats object detection as a set prediction problem. It eliminates the need for specialized components such as anchor boxes and NMS, using transformers to directly predict the final set of detections.
All object detection models are trained using a ResNet50 backbone, except for SSD Lite, which is trained with a MobileNet backbone. The batch size is set to 256 for Faster R-CNN and FCOS and 512 for RetinaNet and SSD Lite. We used a cosine annealing scheduler (Loshchilov & Hutter, 2017) and trained all models for 120 epochs. DETR similarly uses a ResNet50 backbone but has a batch size set to 64. All models are optimized using the AdamW optimizer. For binary object detection model evaluation, we calculate Intersection over Union (IoU) at various thresholds (e.g., IoU0.1, IoU0.3, IoU0.5), which measures how well predicted bounding boxes align with ground truth. IoU is computed by dividing the area of overlap by the area of their union, with higher values indicating better alignment. IoU thresholds define the minimum overlap required for a predicted box to match a ground truth box. (For example, an IoU0.5 threshold means a predicted box must have at least 50% overlap with a ground truth box to be considered a correct detection.) We also assess Mean Average Precision (mAP), including mAP50 and mAP50:95, measuring the model’s precision-recall trade-off and detection accuracy at various IoU thresholds. mAP50 measures precision at an IoU threshold of 0.5, while mAP50:95 averages precision across IoU thresholds from 0.5 to 0.95. Higher mAP scores reflect better detection accuracy and precision.While higher IoU values indicate better accuracy for individual predictions, mAP provides a broader measure of detection performance by capturing precision across different IoU criteria.
5 Results
Our tasks involve identifying a roughly circular well region with a 90m diameter in real life, which translates to less than 30 pixels in satellite imagery due to resizing and other augmentations. This poses a challenge for machine learning models given the heterogeneous nature of the background, including various similarly shaped and sized features of the natural and built environment. Additionally, vegetation can occlude wells in RGB channels, highlighting the importance of near-infrared imagery for guidance. The wells themselves also vary somewhat in shape and can be in various states of disrepair as a result of differing ages and maintenance.
5.1 Binary Segmentation
For the binary segmentation task framing, we train Models (from scratch) using both CNN-based and Transformer-based backbones, considering the prevalent imbalance in the image data due to the small size of wells. Although we did use 3-dimensional, ImageNet initialized weights of the backbone but modified the initial layers afterwards to support 4-dimensional multispectral images. Among our models, as shown in Table 4, the traditional U-Net with EfficientNetB6 backbone performs the best, with CNN-based models showing the highest IOU of $60.4\pm 0.3$ and F1-Score of $64.8\pm 0.4$. While a ResNet50-based backbone achieves the highest Precision of $90.2\pm 2.2$, indicating more accurate predictions of well instances compared to other models. Precision, which reflects the accuracy of our positive detections compared to the ground truth, is crucial. However, a high recall value ensures the model captures most actual well instances, reducing the risk of missing important information. Thus, the Uper-Net model with the highest recall value of $73.1\pm 0.1$, which excels at capturing global context information, appears a good candidate for this task, given a decent precision score. Taking into account both precision and recall, U-Net with EfficientNetB6 backbone performs well, suggesting the utility of a larger backbone with a bigger receptive field. However, ConvNeXt Small provides a trade-off between computational efficiency and performance, requiring fewer GFLOPs and parameters compared to UperNet with Swin Small while achieving similar performance.
5.2 Binary Object Detection
@@
Table 6: Results for object detection task for the FCOS Model with ResNet50 backbone, trained with and without NIR Imagery.
Modality IoU0.1 IoU0.3 IoU0.5 mAP50 mAP50:95
Our evaluation, as shown in Table 5, indicates that while all models perform reasonably well in terms of aligning predicted and actual well locations, performance in the object detection task is overall lower than for segmentation – indicating that potentially segmentation is simply a better framing for this task in real-world settings. The observed gap in performance is likely attributed to the inherent challenge of detecting the small-sized wells within the satellite imagery. It is well-known that single-stage CNN architectures (such as FCOS and SSD) often demonstrate better performance on small object detection than two-stage methods (such as Faster R-CNN), and this aligns with our observations. The exception here is the single-stage method RetinaNet, which, although it has comparable IoU scores, struggles to detect wells accurately. SSD Lite stands out with the highest IoU0.5 score of $65.07\pm 0.03$ and IoU0.3 score of $50.3\pm 0.08$. Whereas all models are quite similar in terms of IoU0.1, the highest score by FasterRCNN is $36.79\pm 1.07$. Thus, SSD Lite and FCOS excel in localization, especially at higher IoU thresholds, while Faster R-CNN is adept at detecting objects with minimal overlap. All models demonstrate low performance in terms of mAP50, which assesses precision-recall trade-off and detection accuracy at an IoU threshold of 0.5. FCOS achieves the of $9.67\pm 1.47$ while SSD Lite achieves a score of $9.76\pm 0.39$. This may be due to these models not producing region proposals confidently enough, especially in instances with a large number of wells. Whereas over a broader evaluation with mAP50:95 which averages precision across IoU thresholds from 0.5 to 0.95. All models apart from RetinaNet provide much better results, with FCOS achieving the highest score of $30.46\pm 3.11$, indicating a decent performance in the identification of well instances. Most notably, the transformer-based model DETR demonstrates the strongest overall performance, achieving strong IoU scores across thresholds and the highest mAP5050{50}50 ($15.22\pm 0.28$) and mAP50:95:5095{50:95}50 : 95 ($38.45\pm 0.32$) scores. These results highlight DETR’s ability to effectively capture global contextual cues essential for accurate well detection.
Table 7: Results for the binary segmentation task for U-Net with ResNet50 backbone, trained with and without NIR imagery.
@@
5.3 Value Added from Dataset Breadth
Benefits of NIR imagery: We find that including the NIR band significantly enhances both object detection and segmentation performance compared to purely RGB data. In segmentation (Table 7), RGB+NIR outperforms RGB in IoU, F1 Score, and Precision while slightly lowering Recall. A similar trend is observed in the object detection (Table 6) with RGB+NIR achieves higher Intersection over Union (IoU) scores across all thresholds (0.1, 0.3, and 0.5) and a considerable increase in mAP@50 and mAP@50:95 . These improvements can be attributed to the enhanced spectral information provided by the NIR band, which is particularly effective in detecting features such as ground depressions that may indicate well sites. These features are often more distinguishable in the NIR spectrum, leading to better performance in both tasks.
Benefits of Using Multiple Well Types. While much past work has focused on active wells, we show that training on such data is insufficient to detect abandoned and suspended wells (which are often significantly harder to e.g. degradation and vegetation growth over time). The results in Table 8 show that including all well types significantly outperforms training on only active wells, improving IoU, F1 score, and recall. Recall in particular is key for monitoring, where missing abandoned wells is more important than erroneously flagging a location as a well.
@@ 6 Limitations
One potential limitation of our work is that we rely on well locations listed by the Alberta Energy Regulator. It is likely that some true well locations are missing in this data, leading to the potential for false negatives in the ground-truth data for this problem. However, it is to be expected that this will not significantly affect the training of algorithms since these labels represent a small fraction of the negative locations in the dataset, and deep learning algorithms are known to be robust to moderate amounts of label noise (see e.g. Rolnick et al. (2017)). Instead the effect may simply be that the reported test accuracy is actually lower than the true value (due to certain correctly predicted well locations being evaluated as false). We hope to investigate such effects further in future work.
@@
While our initial results offer encouragement, a deeper examination reveals a few failure cases arising from visually complex scenarios that highlight key challenges in our detection pipeline (as represented in Figure 8). Our model performs well in typical scenarios, where most patches contain only 1–5 wells. However, in rare high-density regions, performance can degrade —- leading to undercounting or missed detections. Abandoned and suspended wells also pose a challenge due to their subtle visual signatures and minimal surface infrastructure. While near-infrared bands help in some cases, spectral cues alone are not always sufficient. Improving detection in these visually complex and low-signal cases is a key focus for future work.
Figure 8: A few Sample image patches from our dataset presenting qualitative results with predictions of failure cases generated by our Segmentation U-Net and DETR with ResNet50 model.
@@ 7 Conclusion
In this paper, we present the first large-scale benchmark dataset aimed at identifying oil and gas wells, with a focus on abandoned and suspended wells, which are a significant source of greenhouse gases and other pollutants. We combine high-resolution imagery, an extensive database of well locations, and expert verification to create the Alberta Wells Dataset. We frame well identification both in terms of object detection and binary segmentation and evaluate the performance of a wide range of popular deep learning methods on these tasks.
We find that the UperNet Model (with a ConvNexT backbone), in particular, represents the most promising baseline for the binary segmentation task, while for object detection, all models demonstrate more mixed results, with Single Stage Models (such as FCOS and SSD) providing a relatively promising baseline. These results show that the Alberta Wells Dataset represents both a challenging as well as a societally impactful set of tasks.
The value added by the dataset is twofold. First, most global fossil fuel-producing regions do not have databases of well locations comparable to that provided by AER. Alberta’s varied landscape provides an ideal setting for developing algorithms for the detection of wells, which can then be used directly in other locations or fine-tuned. Second, even in Alberta, the list developed by AER is not comprehensive, and many abandoned wells are believed to be missing; remote sensing algorithms can provide candidate locations for domain experts to examine so as to determine the true number and locations of abandoned wells in Alberta.
Overall, we hope that our work will make it possible to (i) better estimate the total effects of well-related methane emissions (currently extremely uncertain (Williams et al., 2020)) and groundwater contamination, (ii) prioritize wells to be plugged and decontaminated. We believe that the scalability of remote sensing will make it an invaluable tool in mitigating the global environmental impact of abandoned oil and gas wells.
Software and Data
Alberta Wells Dataset is available at: https://zenodo.org/records/13743323. Our Benchmarking Code is available at: https://github.com/RolnickLab/Alberta_Wells_Dataset.
Acknowledgments and Disclosure of Funding
This work was funded in part through the Canada CIFAR AI Chairs program and a McGill Tomlinson Scientist Award. We acknowledge support from Planet Labs in the form of their academic subscription program, and are grateful for the help of Tim Elrick in working with Planet data. We also thank Mila and the Digital Research Alliance of Canada for provision of computing facilities, and NVIDIA Corporation for additional computational resources. Special thanks to Francis Pelletier, whose feedback on our data collection pipeline and coding significantly improved our methods.
Impact Statement
@@
Appendix A Hosting, licensing, and maintenance plan
A.1 Hosting & Maintenance
We are hosting the dataset on Zenodo at : https://zenodo.org/records/13743323.
A.2 Data Licensing
The satellite imagery for this project was acquired through Planet Labs’ (PBC, 2024) Education & Research license, which allows the use of the data in publications and the creation of derivative products related to those publications. However, the raw imagery cannot be shared publicly. To adhere to these guidelines, we provide the data in HDF5 format, with the satellite imagery pre-processed to produce a derived product represented as a numpy array from Raster Vector. This process removes all geographic metadata.
This data is for academic use only and should not be used commercially. Proper credit to the current authors, Planet Labs (PBC, 2024), and the Alberta Energy Regulator (AER, 2024) is required when using this data.